Rowing with the current of a river, a rowing team can row 25 mi. in the same amount of time it takes to row 15 mi. against the current. The rate of the rowing team in calm water is 20 mph. Find the rate of the current.
Answer by mananth(16946)

Rowing with the current of a river, a rowing team can row 25 mi. in the same amount of time it takes to row 15 mi. against the current. The rate of the rowing team in calm water is 20 mph. Find the rate of the current.
..
rate of current be x mph
with current rate = x+20 mph
distance = 25 miles
..
time = 25/(20-x)
...
against current speed = x-20 mph
distance = 15
time = 15/(20-x)
..
25/(x+20)= 15/(20-x)
25(20-x)= 15(x+20)
-25x+500= 15x+300
-25x-15x= -500+300
-40x = -200
x= 5 mph
...
m.ananth@hotmail.ca